OpenStudy (anonymous):

PLEASE HELP: solve: (3y-5)^(1/2)=4

OpenStudy (zehanz):

It is:\[\sqrt{3y-5}=4\] You now also know 3y-5=16, because the square root of 16 is 4. Could you do the next step now?

OpenStudy (anonymous):

im not sure wht the next step is

OpenStudy (zehanz):

If 3y-5=16, then what is 3y?

OpenStudy (anonymous):

7

OpenStudy (zehanz):

No...because 7-5 is not 16.

OpenStudy (zehanz):

Oh, I see, you skipped a step ;) 3y=21, so y=7

OpenStudy (zehanz):

Are these multiple choice questions?

OpenStudy (anonymous):

no.

OpenStudy (zehanz):

Then remember it is important to write down the solution step by step!